OpenStudy (anonymous):

Jaime wants to run 15 3/4 miles this week. If he runs 4 3/8 miles a day, how many days will Jaime have to run?

OpenStudy (anonymous):

@amistre64 @gelow @paki

OpenStudy (anonymous):

@TheSmartOne @thomaster

OpenStudy (anonymous):

Someon PLZ help me!!!

OpenStudy (anonymous):

@BlossomCake @lizz123

OpenStudy (panda.exe):

all u have to do is divide them and u will get ur anser

OpenStudy (anonymous):

How can you plz Explain to me?

OpenStudy (panda.exe):

i will give u the sum

OpenStudy (anonymous):

sum of?

OpenStudy (panda.exe):

the anser

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

But then can you explain how to get it plz?

OpenStudy (panda.exe):

he will have to 3.6 days 15 3/4 / 4 3/8 = 63/4 / 35/8 = (63×8) / (4×35) = 504/140 = 18/5 = 3.6

OpenStudy (anonymous):

oops i forgot to give the answer choices Hold On.

OpenStudy (panda.exe):

Ok

OpenStudy (anonymous):

A. 3 1/2 B. 20 1/8 C. 5/18 D. 3 3/5

OpenStudy (anonymous):

The first numbers are whole numbers then its the fraction

OpenStudy (anonymous):

except in C thats just a fraction

OpenStudy (anonymous):

its D Right?

OpenStudy (panda.exe):

c because 18/5 simplafide is 3 3/5

OpenStudy (panda.exe):

I ment d lol

OpenStudy (anonymous):

lol Thx

OpenStudy (anonymous):

Can you help me in more questions plz

OpenStudy (panda.exe):

yes

OpenStudy (anonymous):

ok i am going to post a new question up right now

OpenStudy (anonymous):

I will tag you

OpenStudy (panda.exe):

give medal plz

OpenStudy (anonymous):

i did

OpenStudy (panda.exe):

nvm

OpenStudy (anonymous):

ok

OpenStudy (panda.exe):

ok tag me